OpenStudy (anonymous):

Explain why i^3=-i. What does i^4 equal?

OpenStudy (anonymous):

i^4=i^3*i=-i*i=-i^2=-(-1)=1

OpenStudy (anonymous):

I dont undertand how though

OpenStudy (gw2011):

i^2=-1 (i^2)(i)=(-1)(i)=-i i^4=(i^2)(i^2)=(-1)(-1)=1

OpenStudy (amistre64):

you ever tell time by a clock?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

this is just clock addition .... but it only has 4 hours around it

OpenStudy (amistre64):

i is at 12 oclock; i^2 is at 3oclock; i^3 at 6 oclock; and i^4 at 9 oclock. does that make sense?

OpenStudy (anonymous):

amistre64 u shud try sayin the unit circle...

OpenStudy (amistre64):

unit circle is an advanced concept, but yes :)

OpenStudy (amistre64):

like this

OpenStudy (amistre64):

the reason it changes between is and 1s is that it is actually rotating thru the complex plane;

OpenStudy (anonymous):

amistre64 can u give some of u r medals to me? :D lol

OpenStudy (amistre64):

sure lol

OpenStudy (anonymous):

ya rotation is e^(i\theta)

OpenStudy (amistre64):

theres 1 :)

OpenStudy (anonymous):

thnx...

OpenStudy (amistre64):

if you know how a clock moves from 9 to 3 by adding 6 hours; then this should be easy enough to understand

OpenStudy (amistre64):

its also known as modular addition